OpenStudy (anonymous):

how do I solve this equation? 24x^3y^2 divided by -18x^5y

OpenStudy (mr.math):

This is not an equation.

OpenStudy (anonymous):

you can simplify

OpenStudy (mr.math):

This is an expression that you probably want to simplify.

OpenStudy (across):

\[\frac{24x^3y^2}{-18x^5y}\]can be simplified.

OpenStudy (anonymous):

answer is 4y^2 over -3x^2

OpenStudy (turingtest):

thank you sasdell :\

OpenStudy (anonymous):

your welcome. Why the sad face ?

OpenStudy (turingtest):

because I'm not so sure your helping the student learn.

OpenStudy (anonymous):

fine i'll draw it out

OpenStudy (across):

sasdell, XD, all of us can answer this question in our sleep and within 3 seconds, but where can you see we gave Murg the answer? :p

OpenStudy (anonymous):

ask Turing

OpenStudy (turingtest):

?

OpenStudy (mr.math):

\[{24x^3y^2 \over -18x^5y}=-{4 \over 3}{y^{2-1} \over x^{5-3}}=-{4 y \over 3x^2}.\]

OpenStudy (across):

Murg, the best thing to do would be what Mr.Math did up there. That is, view the fraction as a product of terms\[\frac{24}{-18}\cdot\frac{x^3}{x^5}\cdot\frac{y^2}{y}\]and simplify.

OpenStudy (anonymous):

Ohh, ok, thanks :)

OpenStudy (mr.math):

You're welcome! :)